OpenStudy (anonymous):

How do we simplify: [sqr(2vw^5)][sqr(10v^5w^2)]

OpenStudy (anonymous):

take the w^5 and divide by two as many times, you can do it twice with one left so you get a w^2 outside the sqrt and a w in it then on the second equation do the same with the v, take it out as many times as you can, since its a sqareroot, for every two you take out, you have one on the outside. for the w on the second sqareroot, you have exactly 2, so you can take both out and put one w outside. then simplify by combining like terms

OpenStudy (anonymous):

make sense or do you need me to show step by step?

OpenStudy (anonymous):

so would it be \[w^{2} \sqrt{2vw}\times v ^{2}\sqrt{10vw ^{2}} \] or \[20v ^{6}w ^{7}\]?